# Towards Designing Social Interventions for Online Climate Change Denialism Discussions

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 28th ACM Conference on Computer-Supported Cooperative Work and Social Computing; October 18–22; Bergen, Norwayisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
How can we introduce interventions that motivate conspiracy theory believers to explore scientific inconsistencies and fallacies in conspiracy theories? Building on the previous work observing the role of doubt and deliberation in disengagement from online CT discussion (Phadke et al., 2021; Engel et al., 2023), this paper contributes a theoretically motivated intervention design framework that can be deployed in open discussions on social media. Specifically, we investigate how insider language could be used to contest socially constructed conspiracy theory ideology. Insider language refers to terminologies that people use to discuss conspiracy theories that are unique to the community (Phadke et al., 2021, 2022) For instance, the bolded phrases in the examples above (e.g., “trace gas,” “sea level rise,” “temperature measurements,” etc.) are some specific terms people often use in these communities to describe climate change-related concepts. In this work, we ask:
@@ 3. Method
In the following section, we discuss how we designed, implemented, and deployed our interventions to selected subreddits. We first used the SAGE method to analyze and identify insider language terms specific to the subreddits we are studying. We then designed our interventions and iterated them through qualitative exercises to ensure their effectiveness. Finally, we deployed our interventions in the subreddits in three phases and collected community members’ responses to our post. The entire steps are also shown in Figure 1.
@@ 3.1. Subreddits containing climate denialism
consumer Citing different forms of consumer exploiration as motivation to advocate for climate change theory
@@
Table 1. This table contains examples of the insider terms we used.
@@ 3.2. Insider Language Detection
To ensure that our interventions were relevant and effective in r/climateskeptics, we first investigated how the community discussed the climate change denialism perspective. We first identified the “insider terms” of that community, which refers to the unique language used in the community. Prior research suggests that engaging people with these insider terms can be more effective in promoting critical thinking and challenging conspiracy theory logic (Byford, 2011; Bessi et al., 2015; Lewandowsky et al., 2017; Phadke et al., 2021, 2022).To do so, we used the SAGE (Sparse Additive Generative Models) method, which has been adopted to extract and identify insider language in studying conspiracy theories (Eisenstein et al., 2011; Phadke et al., 2022). SAGE is a method that uses a regularized log-odds ratio to contrast different text corpora’s word distributions, highlighting the words that are uniquely used in each text corpora. In particular, we conducted a comparative analysis of r/climateskeptics against other subreddits (r/conspiracy, r/conservative) and pinpointed terminology distinctive to r/climateskeptics. We excluded words that were generic, that could have other meanings and those that were not frequently used in the last three months. We then conducted open coding on the identified insider language terms (see: Table 1) and highlighted how the community members share their insights and interact with one another. We then analyzed how these terms are used in anti-climate change context and brainstormed counter-arguments to these uses.
@@ 3.5.1. Data
After the final deployment, we collected every response, including a total of 44 follow-up discussions in the thread. In total, we deployed 62 interventions on Reddit (46 were in r/climateskeptics, 16 were in r/conspiracy). 42 of these interventions received responses. 29 of the responses were by the original poster. Figure 3 shows an example intervention and people’s response to it. Additionally, we observed that 36 of 46 interventions deployed in r/climateskeptics were relevant. Some of our interventions misfired because the participants had been discussing different topics than the post that invoked the insider term. Figure 4 shows an example of an out-of-context example. While the original poster mentioned the term “chaos theory,” the discussions in the post was not about this in the context of climate change. Among the 16 interventions deployed in r/conspiracy, 6 were relevant. As we mentioned in the previous section, this phenomenon was more prominent with our interventions deployed in r/conspiracy because we built our interventions using insider language from r/climateskeptics. These terms were slightly different from those of r/conspiracy. We removed the out-of-context interventions from data analysis. In total, we had 42 data points for analyses.
Figure 3. An example of an intervention deployed in r/conspiracy that directly addressed the original poster’s argument.
Figure 4. An example of an intervention in r/consipracy that was out of context because the keywords detected were not associated with the climate change context.
@@ 3.5.2. Analyses
First, we analyzed all the original posts that we intervened on and identified two groups of Redditors that we interacted with. The first group are climate change deniers, who rejected scientific consensus on climate change. The other group are climate change supporters, who supported our post. (Figure 6 & Figure 6). We describe the characteristics of each population in more detail in subsection 4.1. Considering the original posters and all the Redditors who engaged in follow-up discussions, we interacted with 72 Redditors. Among these Redditors, 57 were climate change deniers, and 15 were climate change supporters. Among the 42 relevant interventions deployed, 30 were deployed to climate change deniers and 12 were deployed to pro-climate change users.
@@
Figure 5. This figure shows an example interaction between our intervention and the climate change deniers
Figure 6. This figure shows an example interaction between our intervention and the climate change supporters
Theme Definition
Deniers: evidence Intervention Contained Evidence: Intervention: “I value your perspective. However, have you thought about the current rapid escalation in sea levels? While the mid-holocene saw fluctuations, the pace of change was much slower. Shouldn’t this be a crucial factor in the climate change debate? See here [Linked Resource]” Response: “I’m sorry my friend I do not believe that sea levels are different …stuff still looks just the same as it did when I was a kid.” (D21) Intervention Contained No Evidence: Intervention: “Do you believe that alarmist language around climate change could be a result of the growing body of evidence? Is it possible that these new terms help us better understand the issue?” Response: “There is no body of evidence presented here so the ‘new terms’ only serve to obfuscate the fact.” (D24)
Deniers: original poster’s emotion Original poster was emotional: Intervention: “Your concerns about alarmist narratives are valid, but have you delved into the compelling information offered by climate science? Might this not provide us with a more factual perspective on our changing climate, apart from alarmist views? See here [Linked Resource]” Response: “Evidence of climate change? …Compelling bullshit is what this is.” (D18) Original poster was calm: Intervention: “That’s an interesting perspective. Do you think we should be considering the rate of sea level rise, in addition to the absolute change, when evaluating the impact of climate change? While sea levels during the mid-holocene were different, the rate of change was much more gradual than what we’re experiencing today. See here [Linked Resource]” Response: “The arctic ice melt is further evidence of this. The rate of change is still very slow today. Certainly it’s slow enough to adapt to it calmly.” (D28)
Deniers: responses to intervention with no mention of terminologies “Alarmist” was mentioned: Intervention: “Your caution against alarmist dialogue is noted, but have you explored the breadth of data that climate science offers? Could this not expand our understanding of climate change, leaving alarmist exaggerations aside?” Response: “Alarmist dialogue is wrong. Have you heard of a thing called the fossil record? In which extremes in climate are forever recorded? Let’s include that in any discussion on climate change.” (D27)
Supporters: responding with more rationale Intervention: “It’s fascinating to explore the scientific evidence linking greenhouse gases and the Earth’s temperature, especially in the context of climate change. What’s your take on the topic?” Response: “Infarered energy escapes through water vapor. Water vapor isn’t the issue. It’s co2 and the methane and other harmful gases that us humans have unnaturally put into the atmosphere.” (S12)
Supporters: Responding with more pro-climate change evidence Intervention: “Your input is important! How does the scientific evidence supporting the role of greenhouse gases in the Earth’s temperature impact our discussions about climate change and its potential causes? See this link [Linked resource]” Response: “Understanding greenhouse gases is essential to understanding the current climate crisis. [Linked resource]. Accelerated climate change is threatening the way we go about life …Businesses will have to adapt quickly. (S12)”
@@ 4.1.1. Interventions to Climate Change Deniers
We define climate change deniers as those who oppose the opinions of our intervention and believe in anti-climate change. In particular, deniers expressed strong opinions about climate change scientists and the evidence they provided. For instance, D26 stated that “there is a claim ice is melting” but that “no human on this planet can stand on a beach and say it was here and now it’s here” (Figure 6) to argue against sea level change. This is also a clear marker of climate change deniers.
@@ 4.1.2. Interventions to Climate Change Supporters
We define climate change supporters as those who agree with the climate change perspective. Specifically, this group of people reinforces evidence supporting climate change. For instance, in Figure 6, the original poster clearly states that a set of evidence contributes to global climate change, confirming the supporters’ belief. Generally, the supporters agree with climate change stances, including the rising sea level and greenhouse gas emissions. They also cited evidence to support that stance. For instance, S12 cited external evidence to showcase that “the greenhouse gas emissions are heating up the planet.” Among the 12 interventions deployed to climate change supporters, 3 did not receive a response and 9 received responses from the original posters. The average length of the responses to our interventions was 30 words (Min = 4, Max = 142). The TF-IDF cosine similarity (Christian et al., 2016) between the response and our intervention is 0.52.
@@
In addition, to understand the differences between how climate change deniers and supporters respond to the interventions, we performed a t-test on the sentiment of the responses by the two groups. Our results indicated that there is no significant difference ($t(38)=1.03,p>0.05$). However, this could possibly be due to the small sample size we had, including 30 climate change deniers’ responses and 10 supporters’ responses.
@@ 4.2.1. Responding depending on whether the intervention contains evidence
In contrast, when no evidence is introduced in our intervention, climate change deniers directly asked for more evidence or choose to respond with conspiracy theory supporting evidence. In D2’s response to our intervention without evidence, D2 explicitly said “based on what evidence” and shared that “there are 400,000 glaciers in the world and only a handful have been studied completely.” D2’s response fortified climate change deniers’ own theory and refuted our pro-climate change intervention as “speculation without evidence.” (D2) These examples demonstrate that including evidence in interventions would help avoid deniers’ fixating on the existence of pro-climate change evidence but instead help them focus their attention on having an open debate about climate change.
@@ A.1. Adapting interventions across subreddits
As discussed in subsubsection 3.5.1, when deploying our intervention that originally designed for r/climateskeptics in the community r/conspiracy, some interventions misfired and were out of context (Figure 4). This was because the two communities, though both discussed climate change conspiracy theories, used the same terms in different ways. For instance, as shown in Figure 4, members of r/conspiracy were discussing the use of “chaos theory,” but they did not refer to it in the context of climate change. This mismatch between the terminology and the context it was used in led to our intervention (about climate change) being misfired. Thus, future works exploring intervening in conspiracy theory-related communities should be aware of the uniqueness of each community, and adapt their interventions accordingly. This can help ensure the preciseness and effectiveness of any intervention deployed.

### v2

@@ Abstract.
conspiracy theories, social media
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW251publicationmonth: 11doi: 10.1145/3757432ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Social media
@@ 1. Introduction
How can we introduce interventions that motivate conspiracy theory believers to explore scientific inconsistencies and fallacies in conspiracy theories? Building on the previous work observing the role of doubt and deliberation in disengagement from online CT discussion (Phadke et al., 2021; Engel et al., 2023), this paper contributes a theoretically motivated intervention design framework that can be deployed in open discussions on social media. Specifically, we investigate how insider language could be used to contest socially constructed conspiracy theory ideology. Insider language refers to terminologies that people use to discuss conspiracy theories that are unique to the community (Phadke et al., 2021, 2022). For instance, the bolded phrases in the examples above (e.g., “trace gas,” “sea level rise,” “temperature measurements,” etc.) are some specific terms people often use in these communities to describe climate change-related concepts. In this work, we ask:
@@ 3. Method
In the following section, we discuss how we designed, implemented, and deployed our interventions to selected subreddits. We first used the SAGE (Sparse Additive Generative Models) method to analyze and identify insider language terms specific to the subreddits we are studying (Eisenstein et al., 2011; Phadke et al., 2022). We then designed our interventions and iterated them through qualitative exercises to ensure their effectiveness. Finally, we deployed our interventions in the subreddits in three phases and collected community members’ responses to our post. The entire steps are also shown in Figure 1.
@@ 3.1. Subreddits containing climate denialism
Table 1. This table contains examples of the insider terms we used.
@@
consumer Citing different forms of consumer exploration as motivation to advocate for climate change theory
@@ 3.2. Insider Language Detection
To ensure that our interventions were relevant and effective in r/climateskeptics, we first investigated how the community discussed the climate change denialism perspective. We identified the “insider terms” of that community, which refers to the unique language used in the community. Prior research suggests that engaging people with these insider terms can be more effective in promoting critical thinking and challenging conspiracy theory logic (Byford, 2011; Bessi et al., 2015; Lewandowsky et al., 2017; Phadke et al., 2021, 2022). To do so, we used the SAGE (Sparse Additive Generative Models) method, which has been adopted to extract and identify insider language in studying conspiracy theories (Eisenstein et al., 2011; Phadke et al., 2022). SAGE is a method that uses a regularized log-odds ratio to contrast different text corpora’s word distributions, highlighting the words that are uniquely used in each text corpora. In particular, we conducted a comparative analysis of r/climateskeptics against other subreddits (r/conspiracy, r/conservative) and pinpointed terminology distinctive to r/climateskeptics. We excluded words that were generic, that could have other meanings and those that were not frequently used in the last three months. We then conducted open coding on the identified insider language terms (see: Table 1) and highlighted how the community members share their insights and interact with one another. We then analyzed how these terms are used in anti-climate change context and brainstormed counter-arguments to these uses.
@@ 3.5.1. Data
After the final deployment, we collected every response, including a total of 44 follow-up discussions in the thread. In total, we deployed 62 interventions on Reddit (46 were in r/climateskeptics, 16 were in r/conspiracy). 42 of these interventions received responses. 29 of the responses were by the original poster. 3(a) shows an example intervention and people’s response to it. Additionally, we observed that 36 of 46 interventions deployed in r/climateskeptics were relevant. Some of our interventions misfired because the participants had been discussing different topics than the post that invoked the insider term. 3(b) shows an example of an out-of-context example. While the original poster mentioned the term “chaos theory,” the discussions in the post was not about this in the context of climate change. Among the 16 interventions deployed in r/conspiracy, 6 were relevant. As we mentioned in the previous section, this phenomenon was more prominent with our interventions deployed in r/conspiracy because we built our interventions using insider language from r/climateskeptics. These terms were slightly different from those of r/conspiracy. We removed the out-of-context interventions from data analysis. In total, we had 42 data points for analyses.
(a) An example of an intervention deployed in r/conspiracy that directly addressed the original poster’s argument.
(b) An example of an intervention in r/consipracy that was out of context because the keywords detected were not associated with the climate change context.
(c) An example interaction between our intervention and the climate change deniers.
(d) An example interaction between our intervention and the climate change supporters.
Figure 3. This figure shows examples of how our intervention was deployed and how community members (including both climate change deniers and supporters interacted with our intervention.
@@ 3.5.2. Analyses
First, we analyzed all the original posts that we intervened on and identified two groups of Redditors that we interacted with. The first group are climate change deniers, who rejected scientific consensus on climate change. The other group are climate change supporters, who supported our post. (3(c) & 3(d)). We describe the characteristics of each population in more detail in subsection 4.1. Considering the original posters and all the Redditors who engaged in follow-up discussions, we interacted with 72 Redditors. Among these Redditors, 57 were climate change deniers, and 15 were climate change supporters. Among the 42 relevant interventions deployed, 30 were deployed to climate change deniers and 12 were deployed to pro-climate change users.
@@
Stakeholder Theme Definition and Examples
Deniers Evidence Intervention Contained Evidence: Intervention: “I value your perspective. However, have you thought about the current rapid escalation in sea levels? While the mid-holocene saw fluctuations, the pace of change was much slower. Shouldn’t this be a crucial factor in the climate change debate? See here [Linked Resource]” Response: “I’m sorry my friend I do not believe that sea levels are different …stuff still looks just the same as it did when I was a kid.” (D21) Intervention Contained No Evidence: Intervention: “Do you believe that alarmist language around climate change could be a result of the growing body of evidence? Is it possible that these new terms help us better understand the issue?” Response: “There is no body of evidence presented here so the ‘new terms’ only serve to obfuscate the fact.” (D24)
Original Poster’s Emotion Original Poster was Emotional: Intervention: “Your concerns about alarmist narratives are valid, but have you delved into the compelling information offered by climate science? Might this not provide us with a more factual perspective on our changing climate, apart from alarmist views? See here [Linked Resource]” Response: “Evidence of climate change? …Compelling bullshit is what this is.” (D18) Original Poster was Calm: Intervention: “That’s an interesting perspective. Do you think we should be considering the rate of sea level rise, in addition to the absolute change, when evaluating the impact of climate change? While sea levels during the mid-holocene were different, the rate of change was much more gradual than what we’re experiencing today. See here [Linked Resource]” Response: “The arctic ice melt is further evidence of this. The rate of change is still very slow today. Certainly it’s slow enough to adapt to it calmly.” (D28)
Responses to Intervention with No Mention of Terminologies “Alarmist” was Mentioned: Intervention: “Your caution against alarmist dialogue is noted, but have you explored the breadth of data that climate science offers? Could this not expand our understanding of climate change, leaving alarmist exaggerations aside?” Response: “Alarmist dialogue is wrong. Have you heard of a thing called the fossil record? In which extremes in climate are forever recorded? Let’s include that in any discussion on climate change.” (D27)
Supporters Responding with More Rationale Intervention: “It’s fascinating to explore the scientific evidence linking greenhouse gases and the Earth’s temperature, especially in the context of climate change. What’s your take on the topic?” Response: “Infarered energy escapes through water vapor. Water vapor isn’t the issue. It’s co2 and the methane and other harmful gases that us humans have unnaturally put into the atmosphere.” (S12)
Responding with More Pro-climate Change Evidence Intervention: “Your input is important! How does the scientific evidence supporting the role of greenhouse gases in the Earth’s temperature impact our discussions about climate change and its potential causes? See this link [Linked resource]” Response: “Understanding greenhouse gases is essential to understanding the current climate crisis. [Linked resource]. Accelerated climate change is threatening the way we go about life …Businesses will have to adapt quickly.” (S12)
@@ 4.1.1. Interventions to Climate Change Deniers
We define climate change deniers as those who oppose the opinions of our intervention and believe in anti-climate change. In particular, deniers expressed strong opinions about climate change scientists and the evidence they provided. For instance, D26 stated that “there is a claim ice is melting” but that “no human on this planet can stand on a beach and say it was here and now it’s here” (3(c)) to argue against sea level change. This is also a clear marker of climate change deniers.
@@ 4.1.2. Interventions to Climate Change Supporters
We define climate change supporters as those who agree with the climate change perspective. Specifically, this group of people reinforces evidence supporting climate change. For instance, in 3(d), the original poster clearly states that a set of evidence contributes to global climate change, confirming the supporters’ belief. Generally, the supporters agree with climate change stances, including the rising sea level and greenhouse gas emissions. They also cited evidence to support that stance. For instance, S12 cited external evidence to showcase that “the greenhouse gas emissions are heating up the planet.” Among the 12 interventions deployed to climate change supporters, 3 did not receive a response and 9 received responses from the original posters. The average length of the responses to our interventions was 30 words (Min = 4, Max = 142). The TF-IDF cosine similarity (Christian et al., 2016) between the response and our intervention is 0.52.
@@
In addition, to understand the differences between how climate change deniers and supporters responded to the interventions, we performed a t-test on the sentiment of the responses by the two groups. Our results indicated that there is no significant difference ($t(38)=1.03,p>0.05$). However, this could possibly be due to the small sample size we had, including 30 climate change deniers’ responses and 10 supporters’ responses.
@@ 4.2.1. Responding depending on whether the intervention contains evidence
In contrast, when no evidence is introduced in our intervention, climate change deniers directly asked for more evidence or choose to respond with conspiracy theory supporting evidence. In D2’s response to our intervention without evidence, D2 explicitly said “based on what evidence” and shared that “there are 400,000 glaciers in the world and only a handful have been studied completely.” D2’s response fortified climate change deniers’ own theory and refuted our pro-climate change intervention as “speculation without evidence.” (D2) These examples demonstrate that including evidence in interventions would help avoid deniers’ fixation on the existence of pro-climate change evidence but instead help them focus their attention on having an open debate about climate change.
@@ A.1. Adapting interventions across subreddits
As discussed in subsubsection 3.5.1, when deploying our intervention that originally designed for r/climateskeptics in the community r/conspiracy, some interventions misfired and were out of context (3(b)). This was because the two communities, though both discussed climate change conspiracy theories, used the same terms in different ways. For instance, as shown in 3(b), members of r/conspiracy were discussing the use of “chaos theory,” but they did not refer to it in the context of climate change. This mismatch between the terminology and the context it was used in led to our intervention (about climate change) being misfired. Thus, future works exploring intervening in conspiracy theory-related communities should be aware of the uniqueness of each community, and adapt their interventions accordingly. This can help ensure the preciseness and effectiveness of any intervention deployed.